OpenStudy (anonymous):

Evaluate the given integral using the substitution (or method) indicated. (x + 7)e^(x + 7)^(2) dx; u = (x + 7)^(2)

OpenStudy (anonymous):

int(X^2*e^(-x))dx Substitute s=-x ds=-1dx = - int(e^s*u^2) ds Integrate by parts u=s^2 dv=e^s ds du=2s ds v=e^s = int 2e^s * s ds - e^s * s^2 Factor out the 2: 2 int [ e^s * s ds - e^s s^2 Integrate by parts u=s dv=e^s ds du=1 ds v=e^s becomes = -e^s * s^2 * s - 2 int s^2 ds The integral of e^s is e^s becomes: = - e^s * s^2 + 2 e^s * s - 2 e^s + C Substitute back in s=-x = - e^(-s) * x^2 - 2 e^(-s) * x - 2 e^(-x) + C

OpenStudy (anonymous):

S isn't a part of it, it's just the symbol i couldn't copy and paste